The weight M of an object on Mars varies directly as its weight E on Earth. A person who weighs 95 lb on Earth weighs 38 lb on Mars. How much would a 100-lb person weigh on Mars?
Found 2 solutions by rcmcc, stanbon:
Answer by rcmcc(152)

This is a simple ratio question, because you are looking for the ratio beween E and M you simply have to divide the two numbers.
38/95=0.4
Now that you know the ratio is .04 simply multiply it with the weight of the person on earth
0.4x100=40
A 100 lb person would weigh 40 pounds on mars.

Answer by stanbon(57377)

The weight M of an object on Mars varies directly as its weight E on Earth. A person who weighs 95 lb on Earth weighs 38 lb on Mars.
----------------------------
M = kE
---------
Find "k":
38 = k*95
k= (38/95)
-----
Equation:
M = (38/95)E
----------------------------------
How much would a 100-lb person weigh on Mars?
M = (38/95)*100
M = 40 lbs
---------------------------
Cheers,
Stan H.